OpenStudy (anonymous):

What is the directrix of the parabola given by the equation y = 1/44 x2? (1 point)

OpenStudy (anonymous):

help please

OpenStudy (amistre64):

one method is to equate this to the line y = 1/2x

OpenStudy (anonymous):

what do you mean @amistre64 ?

OpenStudy (amistre64):

 1/44 x^2 = 1/2 x 1/22 x^2 = x 1/22 x = 1; x=22 2a = 22, when a=11